OpenStudy (anonymous):

simplify. -9/5 + (-3) + (-6/5)

OpenStudy (anonymous):

-9/5-3-6/5

OpenStudy (anonymous):

can you simplify that? its the same as: -1(9/5+3+6/5)

OpenStudy (anonymous):

1/3 + 2 ?

OpenStudy (anonymous):

nope, look at it this way: 9/5+6/5 = 15/5 since they have same denominator; 15/5 = 3 , now add the other 3 and get 3+3 =6 multiply this by -1 and get?

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

right

OpenStudy (anonymous):

how would i simplify [a+(-b)] + [b + (-a)]